OpenStudy (anonymous):

What is a possible value for the missing term of the geometric sequence? 1, 250,_, 50,...

OpenStudy (anonymous):

@agent0smith

OpenStudy (***[isuru]***):

let's assume that the missing term is "x" then if the common ration is "r" then 250*r = x also x*r = 50 r = 50/x substitute it in the first one 250*(50/x) = x x^2 = 1250 = nearly .. 111.8 got it :) ?

OpenStudy (***[isuru]***):

srry the last part should be corrected as this x^2 = 12500 x = 111.8

OpenStudy (anonymous):

why does it go up and back down? I'm confused by this question

OpenStudy (anonymous):

@***[ISURU]***

OpenStudy (anonymous):

111.8 isn't one of the answer choices... these are the choices: 1200 650 250 125